A ball is thrown vertically upward with a speed of +18 m/s. (assume g=10 m/s2

)


a. How high does it rise?


1
Expert's answer
2021-09-13T12:15:41-0400

From third equation of motion ,"v^ 2 =u ^2\u22122gh"


"0=u ^2
 −2g\frac{h}{2}"


"h = \frac{u^2}{g}"


= "\frac{18^2}{10}"

h= 32.4m

Therefore, ball rises to 32.4m